A particle undergoes two displacements. The first has a magnitude of 164 cm and makes an angle of 112.8 ◦ with the positive x axis. The resultant of the two displacements is 110 cm directed at an angle of 31.4 ◦ to the positive x axis. Find the magnitude of the second displacement. Answer in units of cm
Answer #1

using vectors.
To add or subtract vectors you add or subtract their components. Vector components are:
x = r * cos(th) y = r * sin(th)
In your problem:
v1 = <x1,y1> v3 = v1 + v2 = <x3,y3>
find v2 = <x2,y2> = v3 - v1
x1 = 16.4m * cos(112.8deg) = -6.35m
y1 = 16.4m * sin(112.8deg) = 15.11m
x3 = 11.0m * cos(31.4deg) = 9.38m
y3 = 11.0m * sin(31.4deg) = 5.73m
x2 = x3 - x1 = 9.38m - 6.35m = 3.03m
y2 = y3 - y1 = 5.73m - 15.11m = -9.38m
Magnitude of a vector is:

Magnitude of 2nd displacement =
| v | = sq root(x^2 + y^2) = ((3.03m)^2 + (9.38m)^2)^0.5 = 9.85m